2/3+1/3(2 1/2+1/2)square



2 3/8 + 1/2(1/3+5/3)exponent of3

Answer by rfer(16322)

2/3+1/3(2 1/2+1/2)^2
2/3+1/3(3)^2
2/3+3
2/3+9/3=11/3=3 1/3
--------------------
2 3/8+1/2(1/3+5/3)^3
2 3/8+1/2(6/3)^3
2 3/8+1/2*27
2 3/8+13 1/2
51/24+324/24
375/24=15 5/8